The radius of a quarter circle is 1 millimeter. What is the quarter circle's perimeter?


Use 3.14 for ​𝜋.


1
Expert's answer
2022-06-12T15:24:40-0400

The perimeter of a quarter circle is equal to a fourth of the circumference of the circle, plus twice the radius


"P=\dfrac{2\pi r}{4}+2r=(\dfrac{\pi}{2}+2)r"

"=(\dfrac{3.14}{2}+2)(1)=3.57(mm)"